The Crescere corporation, an Italian firm, has purchased an asset in Canada which will generate Euro 1.5 millions in annual depreciation charges (converted at the current spot rate) at the end of each year for 10 years. Assuming inflation in Italy will be 6.42%, inflation in Canada will be 3.2%, and the Italian tax rate is 27.5%. What is the expected real value in today’s currency of the depreciation on taxes at the end of year 5 ?

Select the right answer:

(a) Euro 0.797 million
(b) Euro 0.302 million
(c) Euro 0.400 million
(d) Euro 0.352 million
Answer #1

Depreciation tax shield =Depreciation *Tax rate =1,500,000*27.5% =412500 =0.4125 million

Expected Real value at end of year 5=0.4125*(1+Canada Inflation)^5/(1+Italy Inflation)^5 =0.4125*(1+3.2%)^/(1+6.42%)^5=0.3538 million

Option d is correct option because it is closest one